### v1

# Limits on the Statistical Description of Charged de Sitter Black Holes

Abstract
The thermodynamics of de Sitter black holes is complicated by the presence of two horizons and the absence of a globally defined timelike Killing vector. The standard choice of the Gibbons-Hawking Killing vector is at odds with the interpretation of the surface gravity as an acceleration measured by a physical observer at rest. Focusing on four-dimensional Reissner-Nordström de Sitter black holes we show that this issue can be resolved by adopting a normalization originally proposed by Bousso and Hawking, which defines thermodynamic quantities relative to the unique freely-falling observer at a fixed radial coordinate. Within this framework, we derive new first laws for the black hole and cosmological horizon and re-examine the black hole’s heat capacity. We find that the heat capacity remains finite in the near-extremal Nariai limit, thus averting a breakdown of the semi-classical thermodynamic description. However, the heat capacity does vanish in the cold limit, as expected, and for Nariai black holes in the ultracold limit, indicating that fundamental limitations on the statistical description persist in these regimes. We discuss the implications of our results for log-$T$ corrections to near-extremal de Sitter black holes.
1 Introduction
The study of black hole thermodynamics has provided deep insights into the nature of quantum gravity, suggesting a profound connection between spacetime geometry and the microscopic degrees of freedom that make up the black hole. The extension of these concepts to asymptotically de Sitter space introduces a rich and complex structure due to the presence of a cosmological horizon in addition to the black hole horizon(s). This results in several interesting features not present for black holes in asymptotically flat or anti-de Sitter space. For instance, the existence of both black hole and cosmological horizons leads to two notions of temperature. Due to the cosmological horizon, black holes in de Sitter space also have a maximum size.
A primary difficulty in defining the thermodynamics of de Sitter black holes stems from the lack of a globally defined timelike Killing vector. This limitation makes the definition of conserved quantities, such as the total mass or energy, subtle. From the perspective of an observer between the (outer) black hole horizon and the cosmological horizon, it is natural, however, to use a timelike Killing vector in the static patch that coincides locally to a static observer’s local frame. The conventional approach by Gibbons and Hawking uses a normalization of this Killing vector given by $\xi=\partial_{t}$ [1], where $t$ is the standard static time coordinate. In empty de Sitter space, this corresponds to normalizing the Killing vector to 1-1 at the pole of de Sitter space.
[…]
The Bousso-Hawking normalization requires a careful re-examination of the thermodynamical laws for black holes in de Sitter space [^2] The thermodynamics of de Sitter black holes has been previously studied in [5, 6, 2, 7, 8, 9, 3, 4, 10]. Not all thermodynamic properties of black holes depend on the normalization of the Killing vector, but we’ll see it is crucial when computing the heat capacity., as the total energy for example is no longer simply the mass parameter $M$ that appears in the metric in static coordinates [11]. Instead, the physically relevant mass must be defined with respect to this new reference frame. In this paper, we focus on four-dimensional (electrically) charged Reissner-Nordström de Sitter black holes, although our results have a straightforward generalization to higher dimensions and black holes that rotate. We use the Bousso-Hawking normalization to analyze its thermodynamic properties, focusing on the Nariai, (ultra)cold and lukewarm limits. We review the black hole’s characteristic ‘shark fin’ phase diagram, which is bounded by two types of extremal solutions: cold black holes and charged Nariai black holes. These solutions intersect at the ultracold black hole. After that, we show how the thermodynamic laws are modified.
Our key contributions are twofold. First, we derive new first laws for the black hole and cosmological horizons using Bousso-Hawking normalization. In this derivation, we find it convenient to introduce an auxiliary York boundary at $r=r_{\cal O}$, which allows us to derive separate first laws for the black hole and cosmological horizon [12]. In the final result, however, the dependence on the York boundary is absorbed into variations of the other thermodynamic quantities and the first laws relate variations of the observer-normalized mass $\tilde{M}$ to variations in the entropy and charge. Our definition of mass naturally incorporates the redshift factor between the traditional asymptotic mass $M$ and the observer’s reference frame. Using this appropriately modified first law we classify the different thermodynamical regimes in the confined configuration space of charged de Sitter black holes, emphasizing the important role of the lukewarm solutions, corresponding to stable equilibrium states in the canonical ensemble (at fixed charge).
Second, armed with the thermodynamic identities defined using Bousso-Hawking normalization we explore a potential breakdown of thermodynamics in near-extremal limits. In the classic paper [13], it was argued that the statistical description of near-extremal black holes breaks down when the (absolute value of the) heat capacity becomes smaller than one. At this point, the difference between the energy of the black hole (above the ground state) and the extremal energy drops below the typical energy of a Hawking mode, which suggest potentially large (quantum) corrections. This idea was indeed corroborated from the perspective of the Euclidean path integral, which revealed large quantum corrections to the partition function of near-extremal black holes [14, 15, 16, 17, 18, 19, 20]. These corrections modify the entropy to scale logarithmically with $T$, where $T$ is the temperature of the black hole, and are often referred to as log-$T$ corrections. They become large at low temperatures where they lead to a breakdown of the semi-classical approximation. We stress that in this paper we are not computing log-$T$ corrections. Instead, we use the heat capacity as an indicator where the semi-classical description breaks down.
With our refined thermodynamic approach, embracing the consequences of the Bousso-Hawking normalization, we reassess the behavior of the heat capacity of black holes in de Sitter space in the different extremal limits. We show that the heat capacity remains finite and large for a large part of the phase diagram of four-dimensional electrically charged Reissner-Nordström de Sitter black holes. In particular, we do not see a breakdown of the thermodynamic description of near-extremal Nariai black holes sufficiently far away from the ultracold point. This is to be contrasted with the behavior of the heat capacity computed using the normalization by Gibbons and Hawking, which goes to zero in the Nariai limit [10]. We do find that the heat capacity goes to zero in the cold and ultracold Nariai limit, where the scaling of the heat capacity in fact reproduces the one obained using Gibbons-Hawking normalization.
The structure of this paper is as follows. In Sec. 2, we introduce the Reissner-Nordström de Sitter metric and its phase diagram of regular solutions. We then detail the necessity and mechanics of the Bousso-Hawking normalization and derive the associated first laws. Next, we compute the heat capacity in Sec. 3 and discuss the different extremal limits. We derive a list of the near-extremal behavior of the different thermodynamic quantities in Sec. 4 and finally discuss the implications of our findings for the microscopic and statistical understanding of de Sitter black holes in Sec. 5. Some mathematical details are put aside in the appendices.
Note added: while we were completing this paper, we learned of the related work [21] which studies the one-loop gravitational path integral of Reissner-Nordström de Sitter black holes. They focus on the phase of the partition function using the black hole as an observer whereas we study thermodynamics. Our results are complementary.
2 Thermodynamics of Charged de Sitter Black Holes
We first introduce the four-dimensional Reissner-Nördstrom de Sitter black holes. The Einstein-Maxwell action with positive cosmological constant $\Lambda_{4}=3/\ell_{4}^{2}$ is given by
$I=\int\mathrm{d}^{4}x\sqrt{-g}\left[\frac{1}{16\pi G_{4}}\left(R-\frac{6}{\ell_{4}^{2}}\right)-\frac{1}{4}F_{ab}F^{ab}\right]~,$ (2.1)
with $\ell_{4}$ the four-dimensional de Sitter radius. The equations of motion are given by
$R_{ab}-\frac{1}{2}g_{ab}R+\frac{3}{\ell_{4}^{2}}g_{ab}$ $=8\pi G_{4}T_{ab}$ $~,$ (2.2)
$\nabla_{a}F^{ab}$ $=0~.$
The stress tensor is given by
$T_{ab}=F_{ac}F_{b}^{\,\,\,c}-\frac{1}{4}g_{ab}F_{cd}F^{cd}~.$ (2.3)
Focusing on electrically-charged asymptotically de Sitter black holes the relevant metric and gauge field are
$\mathrm{d}s^{2}$ $=-f(r)\mathrm{d}t^{2}+f(r)^{-1}\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega_{2}^{2}~,$ (2.4)
$f(r)$ $=1-\frac{r^{2}}{\ell_{4}^{2}}-\frac{2G_{4}M}{r}+\frac{G_{4}Q^{2}}{4\pi r^{2}}~,$
$A$ $=\left(-\frac{Q}{4\pi r}+\lambda\right)\mathrm{d}t~.$
Here, $\lambda$ is a gauge parameter and the parameters $M$ and $Q$ are typically associated with the mass and charge of the black hole. However, as we will show this identification is somewhat subtle due to the choice of normalization of the timelike Killing vector.
As is well known, see e.g. [22] and references therein, these black hole solution have a phase space of regular solutions that forms a compact ‘shark fin’ (see Fig. 1) bounded by three different notions of extremality. In terms of $(r_{a},r_{b},r_{c})$, i.e. the inner black hole horizon, the outer black hole horizon and the cosmological horizon these extremal solutions are given by
$\text{Cold Black Holes:}\quad r_{a}$ $=r_{b}~,$ (2.5)
$\text{Charged Nariai Black Holes:}\quad r_{b}$ $=r_{c}~,$
$\text{Ultracold Black Hole:}\quad r_{a}$ $=r_{b}=r_{c}~.$
Figure 1: Shark fin-shaped diagram for Reissner-Nordström-de Sitter black holes. The shaded region corresponds to black hole solutions, and the white region to solutions with naked singularities. The upper boundary of the shark fin corresponds to cold black holes $(r_{a}=r_{b})$ and the lower curved boundary to Nariai black holes $(r_{b}=r_{c})$. The two lines meet at the ultracold point $r_{a}=r_{b}=r_{c}$.
Using the relation
$\ell_{4}^{2}=r_{a}^{2}+r_{b}^{2}+r_{c}^{2}+r_{a}r_{b}+r_{b}r_{c}+r_{a}r_{c}~.$ (2.6)
the mass and charge parameters can be expressed the terms of two horizons and $\ell_{4}$ as
$M$ $=\frac{{\cal M}}{G_{4}}=\frac{1}{2G_{4}}\frac{r_{b}^{2}(r_{b}^{2}-\ell_{4}^{2})-r_{c}^{2}(r_{c}^{2}-\ell_{4}^{2})}{(r_{c}-r_{b})\ell_{4}^{2}}~,$ (2.7)
$Q^{2}$ $=\frac{4\pi{\cal Q}^{2}}{G_{4}}=\frac{4\pi}{G_{4}}\frac{r_{c}(r_{c}^{2}-\ell_{4}^{2})-r_{b}(r_{b}^{2}-\ell_{4}^{2})}{(r_{c}^{-1}-r_{b}^{-1})\ell^{2}_{4}}~,$
where we found it convenient to define the rescaled mass and charge $({\cal M},{\cal Q})$.
2.1 Normalization of the Temperature
To define the mass of the spacetime and other thermodynamic quantities we need to specify a timelike Killing vector. In de Sitter space, this is subtle since there exists no global timelike Killing vector. Instead, from the perspective of the observer living inside a cosmological horizon it is natural to use the Killing vector $\xi\propto\partial_{t}$ that is timelike between the outer horizon of the black hole and the cosmological horizon. This still leaves the choice where to normalize the Killing vector to minus one, which we will discuss next.
For a general normalization we will write $\xi=\gamma\partial_{t}$. For asymptotically flat black holes, $\gamma$ is typically chosen to equal one, such that the norm of the Killing vector is normalized to one asymptotically. With this choice, the surface gravity at the black hole horizon has the interpretation of the acceleration needed to keep an object hovering at the horizon, as measured by a freely falling asymptotic observer at a fixed constant radius. Said differently, the locally defined (normalization-independent) acceleration
$a^{\mu}=(\xi^{\nu}\nabla_{\nu}\xi^{\mu})/(\xi^{\rho}\xi_{\rho})~,$ (2.8)
which diverges at the horizon, leads to a finite surface gravity at the black hole horizon $r=r_{h}$ given by
$\kappa=\lim_{r\to r_{h}}\sqrt{(a^{\mu}a_{\mu})(\xi^{\nu}\xi_{\nu})}~,$ (2.9)
due to a diverging redshift factor $\sqrt{-\xi^{\mu}\xi_{\mu}}$ between the horizon and asymptotic infinity.
Applying the same logic to de Sitter space leads to a wrinkle, as there is no analogous asymptotic region in de Sitter space. Nonetheless, Gibbons and Hawking [1] suggested to again pick $\gamma=1$. This is a natural choice for empty de Sitter space, as the location where $\xi^{\mu}\xi_{\mu}=-1$ corresponds to the pole of de Sitter space at $r=0$. In empty de Sitter space, due to the de Sitter isometry group, this corresponds to the stationary coordinates origin of free-falling observers at fixed radius.
This gives the surface gravity of the (observer-dependent) de Sitter horizon an interpretation as the local acceleration times a redshift factor between the freely falling observer and the horizon. As we will show, this interpretation is no longer valid when including a black hole and still selecting $\gamma=1$.
The location where the norm of the Killing vector equals 1-1 is given by $\xi^{\mu}\xi_{\mu}=-\gamma^{2}f(r)=-1$. Then, assuming $r\neq 0$ the quartic polynomial to solve to find this location is
$ar^{4}+br^{3}+cr^{2}+dr+e=0~,$ (2.10)
with
$a$ $=1~,$ (2.11)
$b$ $=0~,$
$c$ $=\ell_{4}^{2}(\gamma^{-2}-1)~,$
$d$ $=(r_{a}+r_{b})(r_{a}+r_{c})(r_{b}+r_{c})~,$
$e$ $=-r_{a}r_{b}r_{c}(r_{a}+r_{b}+r_{c})~.$
The general solution to (2.10) is given in App. A.
Now let us analyze the roots for different choices of normalization. If we take $\gamma=1$, the determinant of this polynomial is negative for $r_{c}\geq r_{b}\geq r_{a}\geq 0$ such that there are only two real roots. For a general charged solution, we find that the positive root lies at a location $r<r_{a}$. In addition, this location corresponds to an accelerating trajectory with $a^{\mu}a_{\mu}\neq 0$. We therefore conclude that when choosing $\gamma=1$, the surface gravity of the black hole is defined with respect to an accelerating observer that lies behind the inner black hole horizon. Obviously, this suggests we use a different, more appropriate, normalization.
In the presence of a (charged) black hole, there is a unique location between the outer black hole and cosmological horizon where an observer is located at a fixed radius, free-falling and stationary with respect to the two horizons. This was pointed out for uncharged black holes in [2] and also applied to charged black holes in [4], where it was referred to as ‘static sphere’ normalization. In this paper, we will refer to this choice as Bousso-Hawking normalization. One can think of this locus as the trajectory where the gravitational attraction of the black hole precisely cancels the cosmological expansion. This special stationary location, fixed between the two horizons, is unstable as free-falling observers at a slightly smaller radius fall through the black hole horizon and observers at a slightly larger radius are expelled through the cosmological horizon. The latter group of observers quickly see the black hole disappear (in a scrambling time) behind the cosmological horizon and settle down in empty de Sitter space. The other group of observers ends up at the black hole singularity. As we are interested in studying black holes states embedded in de Sitter space, it appears reasonable to focus on the unique observer that can probe both horizons. This special radius $r=r_{\cal O}$ is defined by $f^{\prime}(r_{\cal O})=0$, which results in a quartic polynomial with one positive root that lies between the outer black hole horizon and the cosmological horizon. Normalizing the Killing vector to 1-1 at this location means that it coincides locally to the static observer’s worldline, and fixes the normalization factor to $\gamma=1/\sqrt{f(r_{\cal O})}$. Note that this particular choice reduces to the standard normalization in both the empty de Sitter limit and the limit of vanishing cosmological constant ($\ell_{4}\rightarrow\infty$).
Writing the black hole temperature for an arbitrary normalization as $T_{\gamma}=\kappa_{\gamma}/(2\pi)$, we see that the relation between two choices of normalization $\gamma_{1}$ and $\gamma_{2}$ is given by
$T(\gamma_{2})=T(\gamma_{1})\left(\frac{\gamma_{2}}{\gamma_{1}}\right)~.$ (2.12)
This is precisely the expression for the Tolman temperature [23] that expresses a constant temperature $T_{0}$ in terms of the temperature seen by an (accelerating) observer that is redshifted by an amount $\gamma_{1}/\gamma_{2}$.
2.2 Extremal and Near-Horizon Limits
This different choice of normalization is especially important when discussing the thermodynamics of black holes in de Sitter space. As we already alluded to, different choices of normalization lead to different values of the surface gravity at the black hole and cosmological horizon. For an arbitrary normalization, the temperature of the black hole and cosmological horizon are
$T_{b,c}=\frac{\kappa_{b,c}}{2\pi}=\frac{\gamma}{4\pi\ell_{4}^{2}}(r_{c}-r_{b})\left[2+\left(1+\frac{r_{c,b}}{r_{b,c}}\right)^{2}-\left(\frac{\ell_{4}}{r_{b,c}}\right)^{2}\right]~.$ (2.13)
[…]
Indeed, these observations suggest that the finite temperature of de Sitter space is measured by a non-accelerating observer that sits at the pole of the near-horizon dS2 geometry. The Killing vector $\xi$ should therefore be normalized accordingly. To demonstrate this explicitly, let us introduce the following (proper) time and radial coordinates adapted to the freely falling observer.
$\tau$ $=\sqrt{f(r_{\cal O})}\,t~,$ (2.14)
$\rho$ $=\frac{r-r_{\cal O}}{\sqrt{f(r_{\cal O})}}~.$
Expressing the metric (2.4) in these coordinates we obtain
$\mathrm{d}s^{2}=-\left(\frac{f(\rho)}{f(r_{\cal O})}\right)\mathrm{d}\tau^{2}+\left(\frac{f(\rho)}{f(r_{\cal O})}\right)^{-1}\mathrm{d}\rho^{2}+r(\rho)^{2}\mathrm{d}\Omega_{2}^{2}~.$ (2.15)
In general, this metric will look rather complicated. However, to get to a more tame expression we can expand around the Nariai limit. Since the location $r=r_{\cal O}$ is found by solving $f^{\prime}(r)=0$ we expand $f^{\prime}(r)$ around the Nariai limit
$\ell_{4}^{2}r^{3}f^{\prime}(r)=f_{0}+f_{1}(r_{c}-r_{b})+f_{2}(r_{c}-r_{b})^{2}+{\cal O}(r_{c}-r_{b})^{3}~,$ (2.16)
where the coefficients are given by
$f_{0}$ $=-2\left(r-r_{c}\right)\left(-\ell_{4}^{2}r_{c}+r^{2}r_{c}+rr_{c}^{2}+3r_{c}^{3}+r^{3}\right)~,$ (2.17)
$f_{1}$ $=\left(r-2r_{c}\right)\left(\ell_{4}^{2}-6r_{c}^{2}\right)~,$
$f_{2}$ $=-4\left(r-2r_{c}\right)r_{c}~.$
This leads to the following near-horizon Nariai value for $r_{\cal O}$
$r_{\cal O}$ $=r_{c}-\frac{1}{2}(r_{c}-r_{b})+\frac{\ell_{4}^{2}-4r_{c}^{2}}{4r_{c}(6r_{c}^{2}-\ell_{4}^{2})}(r_{c}-r_{b})^{2}+{\cal O}(r_{c}-r_{b})^{3}~,$ (2.18)
$f(r_{\cal O})$ $=\frac{1}{4r_{c}^{2}\ell_{4}^{2}}(r_{c}-r_{b})^{2}(6r_{c}^{2}-\ell_{4}^{2})+{\cal O}(r_{c}-r_{b})^{3}~.$
Note that these expansions are not valid around the ultracold point where $r_{b}=r_{c}=\frac{\ell_{4}}{\sqrt{6}}$.
Plugging this expansion into the metric and taking the limit $(r_{b},r_{c})\to r_{N}$ we find
$\mathrm{d}s^{2}=-\left(1-\frac{\rho^{2}}{\ell_{2}^{2}}\right)\mathrm{d}\tau^{2}+\left(1-\frac{\rho^{2}}{\ell_{2}^{2}}\right)^{-1}\mathrm{d}\rho^{2}+r_{N}^{2}\mathrm{d}\Omega_{2}^{2}~,$ (2.19)
where we defined the two-dimensional de Sitter radius
$\ell_{2}:=\frac{\ell_{4}}{\sqrt{6-\ell_{4}^{2}/r_{N}^{2}}}~.$ (2.20)
[…]
$\tilde{T}_{b}=\frac{\sqrt{6r_{c}^{2}-\ell_{4}^{2}}}{2\pi r_{c}\ell_{4}}=\frac{1}{2\pi\ell_{2}}~.$ (2.21)
indeed corresponds to the temperature measured by a freely falling observer at the pole of the near-horizon dS2 space.
On the cold branch, the temperature goes to zero independent of the choice of normalization, as long the normalization factor $\gamma$ does not diverge in this limit. This is the case for $\gamma=1/\sqrt{f(r_{\cal O})}$.
Another limit of interest is the so-called lukewarm solution, where the black hole and cosmological temperatures coincide. This limit is independent of the choice of normalization and, in terms of the horizon radii, given by $r_{b}+r_{c}=\ell_{4}$. This translates to a charge-to-mass ratio of
$\left(Q/M\right)_{r_{b}+r_{c}=\ell_{4}}=2\sqrt{\pi G_{4}}~,$ (2.22)
which, amusingly, is precisely the charge-to-mass ratio of a four-dimensional extremal Reissner-Nördstrom black hole in flat space. Lukewarm black holes have particularly nice properties, which include the existence of regular Euclidean continuations and stability in the absence of charged decay.
[…]
Figure 2: There is a unique linear trajectory (purple, dashed) that remains within the sharkfin and ends at the ultracold point given by ${\cal Q}=\sqrt{2}{\cal M}-\frac{\ell_{4}}{6\sqrt{3}}$.
In terms of mass and charge parameters this trajectory is given by
$Q=2\sqrt{2\pi G_{4}}M-\ell_{4}\sqrt{\frac{\pi}{27G_{4}}}~.$ (2.23)
Thus, when we are considering variations away from the ultracold point it is necessary to consider variations that obey $\partial Q/\partial M=2\sqrt{2\pi G_{4}}$. [^3] This was already observed in [4, 10].
2.3 Derivation of First Laws with Bousso-Hawking Normalization
It is now of interest to understand how the thermodynamic laws governing the black hole and cosmological horizon depend on the choice of normalization. The precise form of the first law depends on the region under consideration. In [11] the authors derived the first law by considering a Hamiltonian formalism. By integrating conserved quantities from the black hole horizon to either the cosmological horizon or infinity, they obtained different identities.
For instance, integrating from the black hole horizon to the cosmological horizon [11] found the following variational identity
$T_{b}\delta S_{b}+T_{c}\delta S_{c}+\left(\Phi_{b}-\Phi_{c}\right)\delta Q=0~.$ (2.24)
Here, the electric potential is defined as
$\Phi_{b,c}=-\left.\xi^{\mu}A_{\mu}\right|_{r=r_{b,c}}=\gamma\left(\frac{Q}{4\pi r_{b,c}}-\lambda\right)~,$ (2.25)
where we kept the gauge parameter explicit and the entropies are given by the usual expression
$S_{b,c}=\frac{\pi r_{b,c}^{2}}{G_{4}}~.$ (2.26)
The electric charge is given by the integral
$Q=\int_{S^{2}}\star F~.$ (2.27)
Because the normalization of the Killing vector appears as an overall factor in (2.24), it drops out and this identity is valid for an arbitrary normalization.
By considering the region between either the black hole or cosmological horizon and infinity [11] also considered a first law in which the variation $\delta M$ appears, a variation of the mass parameter appearing in the metric (2.4). The appearance of $\delta M$ implies the $\gamma=1$ normalization and the first law is then given by [11]
$\delta M$ $=+T_{b}\delta S_{b}+(\Phi_{b}-\Phi_{\infty})\delta Q~,$ (2.28)
$\delta M$ $=-T_{c}\delta S_{c}+(\Phi_{c}-\Phi_{\infty})\delta Q~.$
We note that the derivation of this first law integrates over a region that goes outside of the (static) region in between the (outer) black hole and cosmological horizon. As such, it is not naturally adapted to the region accessible to a physical observer.
Upon integrating, it can be shown that this first law is compatible with a Smarr formula given by
$M$ $=+2T_{b}S_{b}+(\Phi_{b}-\Phi_{\infty})Q-2P_{\Lambda}V_{b}~,$ (2.29)
$M$ $=-2T_{c}S_{c}+(\Phi_{c}-\Phi_{\infty})Q-2P_{\Lambda}V_{c}~,$
where $P_{\Lambda}=-\frac{\Lambda}{8\pi G_{4}}$ is a pressure term and $V_{b,c}=\frac{4\pi r_{b,c}^{3}}{3}$ is the thermodynamic volume. The derivation of the thermodynamic volume is given in App. B.
Using Bousso-Hawking normalization, one would expect similar first laws and Smarr formulae to hold. To derive the first laws, we find it convenient at an intermediate stage in the derivation to treat the trajectory of the free-falling observer at $r=r_{\cal O}$ as a York boundary. We can then obtain a first law by following the approach of [11] and integrate from the black hole horizon to the location $r=r_{\cal O}$. This implies that we are studying the thermodynamics of the region between the black hole horizon and a freely falling observer. One can study the region between the York boundary and the cosmological horizon in similar fashion by integrating from the York boundary to the cosmological horizon.
One important difference with the standard treatment of a York boundary, however, is the fact that variations of the location $r=r_{\cal O}$ depend on the black hole parameters. We can express (variations of) $r_{\cal O}$ in terms of (variations of) the mass and charge. For this reason, we will only use the formalism of York boundaries as a convenient computational tool but remove its explicit dependence in the final form of the first laws.
As was demonstrated in [12] using the Iyer-Wald formalism, an additional term appears in the first law when adding a York boundary. Writing quantities evaluated using Bousso-Hawking normalization with a tilde, one can obtain the following first laws for the Brown-York energy $E$
$\delta E$ $=+\tilde{T}_{b}\delta S_{b}+(\tilde{\Phi}_{b}-\tilde{\Phi}_{\cal O})\delta Q-\tilde{P}_{\cal O}\delta A_{\cal O}~,$ (2.30)
$\delta E$ $=-\tilde{T}_{c}\delta S_{c}+(\tilde{\Phi}_{c}-\tilde{\Phi}_{\cal O})\delta Q-\tilde{P}_{\cal O}\delta A_{\cal O}~.$
Here $\tilde{P}_{\cal O}$ is a pressure term associated with the York boundary, for the derivation and explicit form see App. B, and $A_{\cal O}=4\pi r_{\cal O}^{2}$ is its area. By subtracting these two first laws one can reproduce the first law for the region between the black hole and cosmological horizon (2.24), which is independent of the normalization. One can now obtain the following Smarr formulae [12]
$E$ $=+2\tilde{T}_{b}S_{b}+(\tilde{\Phi}_{b}-\tilde{\Phi}_{\cal O})Q-2P_{\Lambda}\left(\tilde{V}_{b}-\tilde{V}_{\cal O}\right)-2\tilde{P}_{\cal O}A_{\cal O}~,$ (2.31)
$E$ $=-2\tilde{T}_{c}S_{c}+(\tilde{\Phi}_{c}-\tilde{\Phi}_{\cal O})Q-2P_{\Lambda}\left(\tilde{V}_{c}-\tilde{V}_{\cal O}\right)-2\tilde{P}_{\cal O}A_{\cal O}~.$
Here $\tilde{V}_{r_{b,c,\mathcal{O}}}=\frac{4\pi r_{b,c,\mathcal{O}}^{3}}{3\sqrt{f(r_{\cal O})}}$ is the thermodynamic volume with Bousso-Hawking normalization.
As already mentioned, variations of the (fictitious) York boundary in our case are not independent variations and by comparing (2.30) with (2.28) we see that (using $\Phi_{\infty}=0$ by a suitable choice of gauge)
$\delta E+\tilde{P}_{\cal O}\delta A_{\cal O}+\tilde{\Phi}_{\cal O}\delta Q=\frac{\delta M}{\sqrt{f(r_{\cal O})}}~.$ (2.32)
That is, we can interpret the sum of the Brown-York energy, the pressure term and the electric potential at the York boundary as the redshifted energy variation $\delta\tilde{M}=\delta M/\sqrt{f(r_{\cal O})}$. This removes all explicit dependence on the York boundary and yields our final result for the first laws
$\delta\tilde{M}$ $=+\tilde{T}_{b}\delta S_{b}+\tilde{\Phi}_{b}\delta Q~,$ (2.33)
$\delta\tilde{M}$ $=-\tilde{T}_{c}\delta S_{c}+\tilde{\Phi}_{c}\delta Q~.$
These first laws can, in principle, be integrated to yield an explicit expression for $\tilde{M}$. In practice, this will be difficult in full generality. Luckily, to study the thermodynamic behavior in near-extremal limits we don’t need to know the explicit form of $\tilde{M}$ in full generality; its variation is enough.
Next, we compute the heat capacity of black holes in the different extremal limits. Before we do so, however, let us comment on the behavior of the electric potential.
2.3.1 Finiteness of the Electric Potential
Let us take a closer look at the behavior of the electric potential (2.25) when we apply Bousso-Hawking normalization
[…]
In [10] it was pointed out that since $\sqrt{f(r_{\mathcal{O}})}\sim(r_{b}-r_{c})+\mathcal{O}(r_{b}-r_{c})^{2}$ (see (2.18)) for $|r_{c}-r_{b}|\ll 1$, the above expression diverges in the Nariai limit for the choice of $\tilde{\lambda}=0$.
Despite the electric potential not being directly observable, this could potentially signal a pathology. As we will show here, however, away from the ultracold point this diverging behavior is in fact inherited from subtracting the value of the electric potential at $r\to\infty$ instead of $r=r_{\mathcal{O}}$, translating to $\tilde{\lambda}=0$ or $\tilde{\lambda}=r_{\mathcal{O}}^{-1}$, respectively. Adopting $\tilde{\lambda}=r_{\mathcal{O}}^{-1}$ and using the expressions from (2.18), we arrive at the expressions
$\tilde{\Phi}_{b}$ $=+\frac{\ell_{4}Q}{4\pi r_{b}\sqrt{6r_{b}^{2}-\ell_{4}^{2}}}+{\cal O}(r_{c}-r_{b})~,$ (2.35)
$\tilde{\Phi}_{c}$ $=-\frac{\ell_{4}Q}{4\pi r_{c}\sqrt{6r_{c}^{2}-\ell_{4}^{2}}}+{\cal O}(r_{c}-r_{b})~.$
We note that the electric potentials are finite when taking the Nariai limit, which stems from fully taking into account the observer’s perspective. The finiteness of the potential above is one-to-one related to the finiteness of $\tilde{\Phi}_{b,c}-\tilde{\Phi}_{\mathcal{O}}$ that appeared in the previous section because $\tilde{\Phi}_{b,c}-\tilde{\Phi}_{\mathcal{O}}=\left.\tilde{\Phi}_{b,c}\right|_{\tilde{\lambda}=r_{\cal O}^{-1}}$. In fact, in the Nariai limit we can compactly express the electric potential as
$\lim_{(r_{b},r_{c})\to r_{N}}\tilde{\Phi}_{b,c}=\pm\left(\frac{\ell_{2}}{r_{N}}\right)\frac{Q}{4\pi r_{N}}~,$ (2.36)
where $r_{N}$ is the Nariai radius and $\ell_{2}$ is the two-dimensional radius of the near-horizon de Sitter space. We recognize this as the electric potential for $\gamma=1$ normalization in the Nariai limit times the dimensionless factor $\ell_{2}/r_{N}$ stemming from the redshift factor.
In the ultracold limit $\ell_{2}/r_{N}\to\infty$ and the electric potential (still) diverges, but, as we now explain, from the perspective of the two-dimensional near-horizon de Sitter space this is not surprising. The proper distance between the observer $r=r_{\cal O}$ at the pole of de Sitter space and the black hole horizon is finite and proportional to $\ell_{2}$. In the ultracold limit this proper distance diverges and it is more appropriate to express the electric potential with respect to a different reference point.
Indeed, if we consider the gauge-independent field strength $F=\mathrm{d}A$ expressed in the natural near-horizon coordinates (2.14) in the Nariai limit instead of the electric potential, we find that the electric field is independent of the normalization, finite and constant, as expected.
$F=-\frac{Q}{4\pi r_{N}^{2}}\mathrm{d}t\wedge\mathrm{d}r=-\frac{Q}{4\pi r_{N}^{2}}\mathrm{d}\tau\wedge\mathrm{d}\rho~,$ (2.37)
We conclude that there are no pathological divergences in physical quantities using Bousso-Hawking normalization.
3 Heat Capacity and Breakdown of Thermodynamics
We now have understood how the thermodynamics of black holes in de Sitter space can be described for two different choices of normalization for the timelike Killing vector. In particular, to properly define the thermodynamics using Bousso-Hawking normalization it seems useful to go through an intermediate step and introduce a York boundary at the location of the stationary free-falling observer between the black hole and cosmological horizon. After this, the relevant first laws can be written down in terms of redshifted quantities.
[…]
The picture emerging from these works is that one can think of black holes in de Sitter space as constrained thermal subsystems in a de Sitter heat bath at a temperature given by the cosmological horizon. A similar interpretation was put forward in [3, 4] where it was shown that the probability to nucleate a black hole from a de Sitter heat bath is given by the exponent of minus the total entropy of the system (the sum of the entropy of the black hole and cosmological horizon). The energy needed to nucleate the black hole comes at the expense of the de Sitter heat bath and such a (rare) fluctuation decreases the total entropy of the system.
Introducing a (fictitious) York boundary at the location of the observer at $r=r_{\cal O}$, essentially allowed us to treat the region between the black hole horizon and $r_{\cal O}$ and the region between $r_{\cal O}$ and the cosmological horizon as two separate thermodynamic systems. In fact, [12] used a York boundary to rigorously be able to define the thermodynamic ensemble in de Sitter space and then taking the limit where the York boundary was removed. In our case, the York boundary arises naturally at the $r=r_{\cal O}$ and its contribution is absorbed into the other thermodynamic quantities.
It is of particular interest to study the limitations of this thermodynamic description. In a classic paper by Preskill et al. [13] it was argued that the thermodynamic description of asymptotically flat Reissner-Nördstrom black holes breaks down close to extremality. At low temperatures, the thermal energy drops below the typical energy of a Hawking quantum leading to a situation analogous to a quantum freezeout in statistical systems. The inability to absorb heat due to absence of available degrees of freedom is indicated by a heat capacity at fixed charge that goes to zero in the extremal limit. The point where the statistical description becomes unreliable corresponds to a heat capacity of order one. More recent works (see [27] for a review) have indeed confirmed that the near-extremal statistical description of Reissner-Nördstrom black holes is significantly modified due to large fluctuations around the saddle point in the Euclidean path integral. Taking these fluctuations into account leads to corrections to the entropy (and more generally the partition function) that scale logarithmically with temperature [14, 15, 16, 17, 18, 19, 20].
Before we discuss the charged case, let us first consider the behavior of uncharged near-Nariai black holes. Using the first law for uncharged black holes, given by $\delta\tilde{M}=\tilde{T}_{b}\delta S_{b}$, we can see that the near-Nariai behavior of the mass is given by
$Q=0:\qquad\delta\tilde{M}=-\frac{\pi\ell_{4}^{2}}{G_{4}}\delta\tilde{T}_{b}+{\cal O}(\tilde{T}_{b}-\tilde{T}_{N})^{2}~,$ (3.1)
where $\tilde{T}_{N}=(2\pi\ell_{2})^{-1}$ is the temperature of the black hole in the Nariai limit. The analogous quantity for asymptotically flat Reissner-Nordström black holes is the (change in the) energy above the (extremal) ground state energy. In that case, the change in the energy above extremality can be interpreted as the energy available to produce Hawking modes. In the de Sitter case, we give (3.1) the same interpretation as the energy available to Hawking modes, with the important difference that now there is a maximum energy state, which explains the minus sign.
This near-extremal behavior should be contrasted with the results of [10] for de Sitter black holes, which were obtained using $\gamma=1$ normalization. First, the dependence of the mass on temperature is linear instead of quadratic. Second, the available energy remains finite and large in the Nariai limit. Still assuming that the energy of Hawking modes have a linear dependence on the temperature, this shows that available energy never dips below the thermal energy. Hence, there is no indication that the thermodynamic description breaks down in the (uncharged) Nariai limit.
This conclusion can also be reached by examining the heat capacity of the black hole system. Using the first law, this heat capacity can be expressed as
$\tilde{C}_{b}=\tilde{T}_{b}\left(\frac{\partial S_{b}}{\partial\tilde{T}_{b}}\right)=\frac{\partial\tilde{M}_{b}}{\partial\tilde{T}_{b}}~.$ (3.2)
Using (3.1) we see that the heat capacity in the (uncharged) Nariai limit does not go to zero, but is large and given by
$\tilde{C}_{b}=-\frac{\pi\ell_{4}^{2}}{G_{4}}~,$ (3.3)
where the minus sign simply reflects the fact that the available energy is an energy below the maximum Nariai mass.
The situation will be different when we include charge, as the temperature $\tilde{T}_{b}$ of Nariai black holes becomes smaller when increasing the charge. To assess the energy below the Nariai mass it is then natural to work at fixed charge. [^4] As already mentioned, one needs to be careful near the ultracold point where variations at fixed charge exit the sharkfin. We therefore now turn our attention to computing heat capacities for the different extremal limits.
3.1 Heat Capacity at Fixed Charge
Using the $\gamma=1$ normalization, it is easily demonstrated that the heat capacity of Reissner-Nördstrom de Sitter black holes also vanishes in the Nariai limit, see e.g. [10, 28]. One should not jump to the conclusion, however, that Nariai black holes have a similar breakdown in their statistical description. Using Bousso-Hawking normalization, the heat capacity is given by
$\tilde{C}_{b}^{Q}=\tilde{T}_{b}\left(\frac{\partial S_{b}}{\partial\tilde{T}_{b}}\right)_{Q}~,$ (3.4)
where $Q$ denotes that the charge is kept fixed. In Fig. 3 we display the heat capacity as a function of the black hole radius. For comparison, we also include the heat capacity defined using $\gamma=1$ normalization
Figure 3: Heat capacity of an uncharged black hole in de Sitter space for $\gamma=1$ normalization and Bousso-Hawking normalization as a function of radius. We used $\ell_{4}=G_{4}=1$.
From Fig. 3 we see that the heat capacity $\tilde{C}_{b}$ of an uncharged black hole is monotonically decreasing as a function of black hole radius. In the Nariai limit $r_{b}\to r_{N}=\frac{\ell_{4}}{\sqrt{3}}$, this reduces to
$\text{Uncharged Nariai:}\qquad\lim_{r_{b}\to r_{N}}\tilde{C}_{b}=-\frac{\pi\ell_{4}^{2}}{G_{4}}~,$ (3.5)
which agrees with (3.3). The heat capacity goes to zero for small (Planck mass) black holes, as it should. Concretely, in the semi-classical limit $\sqrt{G_{4}}/\ell_{4}\ll 1$ we find that $\tilde{C}_{b}=-1$ for Planck mass black holes with radius $r_{b}\sim\sqrt{G_{4}}$. Hence, general relativity breaks down in an ‘ordinary’ way for these black holes. We conclude that, expressed in terms of the redshifted energy $\tilde{M}$, in the Nariai limit there are still a large number of degrees of freedom available that can absorb heat and the statistical description does not break down.
This picture changes when we include charge. By increasing the charge, we can lower the temperature of the black hole and we can end up in a situation where the available energy drops below the Hawking temperature. This is demonstrated very clearly for cold black holes which have a vanishing temperature along the entire cold branch. The near-extremal heat capacities are identical in the different normalizations. This can be attributed to the fact that the leading order term in an expansion of $f({r_{\cal O}})$ around the cold point is constant, as we will explain in more detail in Sec. 4. We find that the heat capacity at fixed charge is
$\text{Cold:}\qquad\tilde{C}_{b}^{Q}=C_{b}^{Q}=\frac{\pi r_{b}^{2}}{G_{4}}\left(\frac{r_{b}}{r_{a}}-1\right)+{\cal O}(r_{b}-r_{a})^{2}~.$ (3.6)
The vanishing of the heat capacity in the extremal limit suggests a breakdown of the thermodynamic description, similar to extremal black holes in flat space. Note, however, that in flat space charged black holes thermodynamically evolve to the extremal state by emitting neutral Hawking radiation. In de Sitter space, at fixed charge, black holes evolve towards the maximal entropy lukewarm state. For masses below the lukewarm point, the cosmological temperature is higher than the black hole temperature, causing the black hole to evolve towards the lukewarm line instead of the cold extremal state. In other words, in de Sitter space one would need to do something special (throw in some extra charge or force the black hole to lose some mass) to end up in the non-equilibrium cold extremal state.
Turning to near-Nariai black holes, the heat capacity at fixed charge is given by
$\text{Charged Nariai:}\qquad\tilde{C}_{b}^{Q}$ $=-\frac{\pi r_{b}^{2}}{G_{4}}\left(\frac{6r_{b}^{2}-\ell_{4}^{2}}{4r_{b}^{2}-\ell_{4}^{2}}\right)+{\cal O}(r_{c}-r_{b})~,$ (3.7)
$C_{b}^{Q}$ $=-\frac{\pi r_{b}^{2}}{G_{4}}\left(\frac{r_{c}}{r_{b}}-1\right)+{\cal O}(r_{c}-r_{b})^{2}~.$
This expression reveals some interesting physics. First, we notice that at $r_{b}=\ell_{4}/2$ the heat capacity $\tilde{C}_{b}^{Q}$ diverges. This point corresponds to the intersection of the lukewarm line and the Nariai line at ${\cal Q}/\ell_{4}=1/4$. A divergent heat capacity is indicative of a phase transition [29] (however, see [30]). Indeed, the qualitative behavior of the black hole states change when crossing the lukewarm line. Black holes above the lukewarm line (in Fig. 1) have a temperature that is lower than that of the cosmological horizon. Below the lukewarm line, the heat capacity is negative and above the lukewarm line the heat capacity is positive, see Fig. 4. We will come back to the interpretation of this divergence in Sec. 3.2 and show how working in a different ensemble gives a finite and positive heat capacity.
Figure 4: Heat capacity at fixed charge as a function of radius of charged Nariai black holes in de Sitter space. We used $\ell_{4}=G_{4}=1$.
Second, above the lukewarm line the heat capacity is positive and vanishes in the ultracold point $r_{b}=\ell_{4}/\sqrt{6}$. This suggests that the statistical description breaks down along the Nariai line when $\tilde{C}_{b}^{Q}<1$, which corresponds to
$r_{b}<\frac{\ell_{4}}{\sqrt{6}}\left(1+\frac{G_{4}}{\pi\ell_{4}^{2}}\right)+{\cal O}\left(G_{4}^{2}/\ell_{4}^{4}\right)~,$ (3.8)
i.e. for black holes very close to the ultracold point. Approaching the ultracold point from the Nariai line this translates to a temperature of
$\tilde{T}_{b}>\frac{\sqrt{3G_{4}}}{\ell_{4}^{2}\pi^{3/2}}+{\cal O}\left(G_{4}^{2}/\ell_{4}^{4}\right)~.$ (3.9)
For temperatures smaller than this, we expect the thermodynamic description to break down.
One should be careful, however, when approaching the ultracold point. As explained, for black hole variations away from the ultracold point that remain within the sharkfin there is a unique trajectory. This implies that, when expanding around the ultracold point, one cannot consider variations where the charge is kept fixed. Instead, any variation should obey $\delta Q=2\sqrt{2\pi G_{4}}\delta M$ corresponding to a non-standard thermodynamic ensemble. We will come back to this point, but for now we summarize the different heat capacities at fixed charge in Tab. 1.
$C_{b}^{Q}$ $\tilde{C}^{Q}_{b}$
Nariai: 0 $\frac{\pi r_{b}^{2}}{G_{4}}\left(\frac{6r_{b}^{2}-\ell_{4}^{2}}{4r_{b}^{2}-\ell_{4}^{2}}\right)$
Cold: 0 0
Ultracold: 0 0
Nariai $\cap$ Lukewarm: 0 $\pm\infty$
Table 1: Heat capacity at fixed charge of the different extremal black holes for both $\gamma=1$ and Bousso-Hawking normalization. The sign of the divergence in $\tilde{C}_{b}^{Q}$ at the intersection of the Nariai and lukewarm line depends if we approach that point from above or below.
3.2 Emission of Charged Particles
Up to now, we only studied the heat capacity of black holes at fixed charge. This is appropriate to study the thermodynamic behavior of black holes that only exchange (neutral) heat, but no charge. However, in general we expect the decay to be governed by (a combination of) Schwinger pair production and neutral Hawking radiation. [^5] It is particularly important to understand the possible decay channels of near-Nariai black holes through Schwinger pair production [9, 22]. When this is the case, we need a different ensemble to describe the thermodynamics of the black hole system.
Two cases where we already encountered this were the intersection of the lukewarm and Nariai line and the ultracold point. In the first case, we found that the heat capacity at fixed charge diverges. [^6] As we will show in more detail later, the heat capacity at fixed charge at the lukewarm line is finite in general. The heat capacity only diverges at the intersection of the lukewarm and Nariai line. In the second case, variations away from the ultracold point require variations in both the charge and mass to remain within the sharkfin.
3.2.1 Heat Capacity at Fixed Charge-to-Mass Ratio
As mentioned, the existence of the Schwinger effect necessitates the discussion of black hole thermodynamics in an ensemble where charge can vary. We will be interested in the heat capacity under Bousso-Hawking normalization, but with the constraint $\delta Q/\delta M=z$ when taking variations. [^7] We could also have expressed this constraint in terms of $\delta Q/\delta\tilde{M}$ which would introduce an additional redshift factor. Under this constraint, heat flow through both horizons equals
$\tilde{T}_{b}\delta S_{b}=+(1-z\Phi_{b})\delta\tilde{M}~,$ (3.10)
$\tilde{T}_{c}\delta S_{c}=-(1-z\Phi_{c})\delta\tilde{M}~.$
Note that the combination $z\Phi_{b,c}$ appearing here does not have a tilde. For uncharged radiation ($z=0$) the heat flow through the black hole horizon always equals minus the heat flow through the cosmological horizon. Including charge, this is no longer true. We can for example consider the situation where there is positive (or vanishing) heat flow through both horizons, which occurs when
$\tilde{T}_{b}\delta S_{b}=\tilde{T}_{c}\delta S_{c}\quad\Rightarrow\quad z=\frac{2}{\Phi_{b}+\Phi_{c}}~.$ (3.11)
In the Nariai limit $\Phi_{b}=\Phi_{c}$ and at this point both heat flows vanish when (3.11) is satisfied.
Now remember that at the intersection of the lukewarm and Nariai line, the heat capacity at fixed charge diverged. This implies that, at that point, the heat flow through the black hole horizon is independent of the temperature. To obtain a finite value for the heat capacity, we need to allow for charge exchange such that the heat flow vanishes. Concretely, we can only obtain a finite value for the heat capacity when $z=\Phi_{b}^{-1}$, at which point we find
$r_{b}=r_{c}=\frac{\ell_{4}}{2}:\quad\tilde{C}_{b}^{z}=\tilde{T}_{b}\left(\frac{\partial S_{b}}{\partial\tilde{T}_{b}}\right)_{z=\Phi_{b}^{-1}}=S_{b}~.$ (3.12)
We now also return to the ultracold point. Similar to the case at the intersection of the lukewarm and Nariai line, the value of $z$ for which the right-hand side of (3.10) vanishes is given by $z=\Phi_{b}^{-1}$. In the case of the ultracold black hole this translates to
$r_{b}=r_{c}=\frac{\ell_{4}}{\sqrt{6}}:\quad z=2\sqrt{2\pi G_{4}}~,$ (3.13)
which is precisely the value corresponding to the purple trajectory described in Fig. 2 necessary to consider variations that remain within the sharkfin. We summarize the heat capacities at fixed $z$ in Tab. 2.
$C_{b}^{z}$ $\tilde{C}_{b}^{z}$
Ultracold: 0 0
Nariai $\cap$ Lukewarm: 0 $S_{b}$
Table 2: Heat capacity at fixed charge-to-mass ratio $z=\Phi_{b}^{-1}$ for both $\gamma=1$ and Bousso-Hawking normalization. $S_{b}$ denotes the entropy of the black hole.
3.3 A Detailed Look at the Lukewarm Line
Having understood the different heat capacities relevant for different parts of the phase diagram, we will now take a closer look at lukewarm black holes. As we already noted, this is a regime of interest for various reasons.
Foremost, it is a thermal equilibrium state where the temperatures of the cosmological and black hole horizon are equal. A corollary of this fact is that lukewarm black holes admit regular Euclidean continuations. Furthermore, because lukewarm black holes mark the boundary between parts of the phase diagram where their temperature is larger or smaller than the temperature of the cosmological horizon they play an important role in determining the dynamic evolution in the sharkfin. Specifically, for fixed charge the lukewarm solution corresponds to a maximum (total) entropy state. This means that the semi-classical thermodynamical evolution will never be able to decrease the mass below the lukewarm value to reach the cold limit. This is because below the lukewarm point (at fixed charge), the cosmological horizon will have a higher temperature than the black hole and there is net energy flux into the black hole through Hawking radiation, increasing the black hole mass. In the past, this semi-classical evolution has been studied using Gibbons-Hawking normalization [9, 31]. Here we will study the behavior of black hole states near the lukewarm line using Bousso-Hawking normalization instead.
Lukewarm black holes are defined by
$\begin{cases}f(r_{b})=f(r_{a})=0\\ f^{\prime}(r_{b})=-f^{\prime}(r_{c})\end{cases},$ (3.14)
where the first equation defines the horizons and the second relates the temperature as
$\tilde{T}_{b,c}=\frac{|f^{\prime}(r_{b,c})|}{4\pi\sqrt{f(r_{\cal O})}}~.$ (3.15)
The metric function thus has the following decomposition
$f(r)=-\frac{1}{(r_{b}+r_{c})^{2}r^{2}}(r-r_{b})(r-r_{c})\left[r^{2}+(r_{b}+r_{c})r-r_{b}r_{c}\right]~,$ (3.16)
from which one can read off the parameters
$\ell_{4}$ $=r_{b}+r_{c}~,$ (3.17)
$\mathcal{M}$ $=\frac{r_{b}r_{c}}{\ell_{4}}~,$
$\mathcal{Q}^{2}$ $=\frac{r_{b}^{2}r_{c}^{2}}{\ell_{4}^{2}}~,$
i.e., lukewarm black holes satisfy the condition $\mathcal{M}=|\mathcal{Q}|$.
Because the Hawking temperature of the black hole and the cosmological horizon is the same, one would expect the lukewarm black holes to be in equilibrium when the spacetime is filled with Hawking radiation. This equilibrium state will be broken, however, if the Schwinger effect can happen, subsequently putting the system in an ensemble satisfying $\delta\mathcal{Q}/\delta\mathcal{M}=\mathcal{Z}$, where the $\mathcal{Z}$ is determined by the charged particle in question. [^8] The fixed charge-to-mass ratio constraint can be written in terms of $z=\delta Q/\delta M$ or $\mathcal{Z}=\delta{\cal Q}/\delta{\cal M}$. We’ll use them interchangeably.
In fact, the Schwinger effect always happen for lukewarm black holes for any charged particle spectrum, unlike cold black holes with a threshold value for the charge-to-mass ratio. The presence of a threshold charge-to-mass ratio for the Schwinger effect to take place has been used as a motivation for the Weak Gravity Conjecture [32, 33]. The absence of a threshold value can be seen from the near-horizon structure of the lukewarm black hole.
We expand the metric function near the black hole as
$f(r)=C\rho+{\cal O}(\rho^{2}),$ (3.18)
where $\rho=r-r_{b}$ and $C=f^{\prime}(r_{b})=\frac{2(r_{c}-r_{b})}{(r_{b}+r_{c})^{2}}$. The metric to leading order is
$ds^{2}$ $=-C\rho\,\mathrm{d}t^{2}+\frac{\mathrm{d}\rho^{2}}{C\rho}+(r_{b}+\rho)^{2}\mathrm{d}\Omega_{2}^{2}$ (3.19)
$\simeq-\alpha^{2}x^{2}\mathrm{d}t^{2}+\mathrm{d}x^{2}+r_{b}^{2}\mathrm{d}\Omega_{2}^{2}~,$
where we made the change of coordinates $\rho=\frac{C}{4}x^{2}$ and defined the acceleration $\alpha=\frac{C}{2}$. The above metric describes the local ${\rm Rindler}_{2}\times S^{2}$ space. Similarly, the near-horizon geometry of the cosmological horizon can also be identified to be ${\rm Rindler}\times S^{2}$ with a different $S^{2}$ size.
With the coordinate expansion, the gauge field near the black hole horizon can be expressed as
$A_{t}=-\frac{Q}{4\pi r}$ $\simeq A_{t}^{(0)}+\frac{Q}{4\pi r_{b}}\frac{\rho}{r_{b}}$ (3.20)
$=A_{t}^{(0)}+\frac{Q}{4\pi}\frac{C}{4r_{b}^{2}}x^{2},$
where $A_{t}^{(0)}$ is a constant piece. Since $F=\mathrm{d}A$ is proportional to the two-dimensional volume form $\sqrt{-g_{2}}\mathrm{d}^{2}x$, there is a constant electric field $E=\frac{Q}{4\pi r_{b}^{2}}$ in the near-horizon two-dimensional Rindler space. The same holds true at the cosmological horizon again with a weaker electric field strength.
The Schwinger production rate is non-zero in the presence of charged fields, as can be inferred from the isometry between Rindler space and Minkowski space. In fact, the instantons that give rise to the Schwinger rate are isometric to those in Minkowski space. This is true because the solutions to the equations of motion are preserved under coordinate transforms. The ‘classical’ part of the Schwinger production rate is invariant, thus only the coefficient related to the quantum fluctuations, sensitive to the choice of vacua, is modified. The total production rate therefore takes the form of the Schwinger rate in Minkowski space, rescaled to take into account the acceleration of the Rindler observer.
Notably, the above implies that, unlike extremal black holes whose Schwinger effect see a threshold value in the charge-to-mass ratio of the produced particle, lukewarm black holes emit charged radiation regardless of the particle spectrum. Depending on the charge-to-mass ratio of the particle, the Schwinger effect can either take the black hole above or below the lukewarm line, where Hawking radiation then takes action and drives the system back to the lukewarm solution.
Returning to the thermodynamics of lukewarm black holes, we will compute the heat capacity subject to the Schwinger effect. Using the general expression of the mass and charge parameters in terms of the black hole and cosmological horizons, the general variations are
$\begin{cases}\delta\mathcal{M}=\frac{r_{b}^{2}(r_{b}^{2}-\ell_{4}^{2})-r_{c}^{2}(r_{c}^{2}-\ell_{4}^{2})+2(2r_{b}^{2}-\ell_{4}^{2})r_{b}(r_{c}-r_{b})}{2\ell_{4}^{2}(r_{c}-r_{b})^{2}}\delta r_{b}-\frac{r_{b}^{2}(r_{b}^{2}-\ell_{4}^{2})-r_{c}^{2}(r_{c}^{2}-\ell_{4}^{2})+2(2r_{c}^{2}-\ell_{4}^{2})r_{c}(r_{c}-r_{b})}{2\ell_{4}^{2}(r_{c}-r_{b})^{2}}\delta r_{c}\\ \delta(\mathcal{Q}^{2})=\frac{1}{(r_{c}^{-1}-r_{b}^{-1})^{2}\ell_{4}^{2}}\left\{\left[(-3r_{b}^{2}+\ell_{4}^{2})(r_{c}^{-1}-r_{b}^{-1})-\frac{r_{c}(r_{c}^{2}-\ell_{4}^{2})-r_{b}(r_{b}^{2}-\ell_{4}^{2})}{r_{b}^{2}}\right]\delta r_{b}\right.\\ \left.-\left[(-3r_{c}^{2}+\ell_{4}^{2})(r_{c}^{-1}-r_{b}^{-1})-\frac{r_{c}(r_{c}^{2}-\ell_{4}^{2})-r_{b}(r_{b}^{2}-\ell_{4}^{2})}{r_{c}^{2}}\right]\delta r_{c}\right\}\end{cases}.$ (3.21)
Restricting to the lukewarm line we obtain
$\begin{cases}\delta\mathcal{M}=-\frac{1}{\ell_{4}^{2}}(r_{b}^{2}\delta r_{b}+r_{c}^{2}\delta r_{c})~,\\ \delta\mathcal{Q}=-\frac{1}{\ell_{4}}(r_{b}\delta r_{b}+r_{c}\delta r_{c})~.\end{cases}$ (3.22)
With the above expression, we can rewrite $\delta\mathcal{Q}=\mathcal{Z}\delta\mathcal{M}$ as
$\delta r_{c}=h(r_{b},r_{c})\delta r_{b}\equiv-\frac{r_{b}}{r_{c}}\frac{1-\mathcal{Z}\frac{r_{b}}{\ell_{4}}}{1-\mathcal{Z}\frac{r_{c}}{\ell_{4}}}\delta r_{b}.$ (3.23)
We next compute the heat capacity according to
$\tilde{C}_{b}^{z}=\tilde{T}_{b}\left.\frac{\delta S_{b}}{\delta\tilde{T}_{b}}\right|_{z}~,$ (3.24)
where the charge-to-mass ratio specifies the ensemble. We note that $S_{b}=\pi r_{b}^{2}/G_{4}$ and
$\tilde{T}_{b}=\frac{f^{\prime}(r_{b};r_{b},r_{c})}{4\pi\sqrt{f(r_{\cal O};r_{b},r_{c})}}~.$ (3.25)
The semi-colon separates the variable $r$, from the parameters $r_{b,c}$. Later on, the parameter dependence will be left implicit. Derivatives of the radial coordinate will be denoted as a prime whereas derivatives with respect to the parameters will be denoted as partial derivatives. Further, when evaluating the metric function at a particular location, we will omit the argument and indicate this with a subscript. For instance, $f^{\prime}_{\cal O}\equiv\frac{d}{dr}\left.f(r;r_{b},r_{c})\right|_{r=r_{\cal O}}$ is a function of $r_{b,c}$. Inserting these definitions we obtain
$\tilde{C}_{b}^{z}=\frac{2\pi r_{b}f^{\prime}_{b}}{\sqrt{f_{\cal O}}}\left[\frac{\delta}{\delta r_{b}}\left(\frac{f^{\prime}_{b}}{\sqrt{f_{\cal O}}}\right)\right]^{-1}~.$ (3.26)
While the quantities evaluated at $r=r_{b}$ can be simplified after restricting to the lukewarm line, the heat capacity in Bousso-Hawking normalization generally requires numerical evaluation due to the observer location $r_{\cal O}$ being the root of a quartic equation, which has a lengthy expression in general. The behavior of $\tilde{C}^{z}_{b}$ with different choices of charge-to-mass ratio is shown in Fig. 5.
Figure 5: Heat capacity of lukewarm black holes in various ensembles specified by $\mathcal{Z}$. From top to bottom, the three curves correspond to $\mathcal{Z}=2.2,2,1.8$ respectively. When $\mathcal{Z}>2$, $\tilde{C}^{z}_{b}\rightarrow\infty$ at the lukewarm-Nariai intersection point where as for $\mathcal{Z}<2$, $\tilde{C}^{z}_{b}\rightarrow$-$\infty$. For the later case, $\tilde{C}^{z}_{b}$ has a zero at the point $1-\frac{\mathcal{Z}r_{c}}{\ell_{4}}=0$ which corresponds to $\delta S_{b}=0$.
From (3.10), we saw that $\delta S_{b}=0$ when $z=\Phi_{b}^{-1}$ Along the lukewarm line, which corresponds to $z=2\sqrt{\pi G_{4}}\ell_{4}r_{c}^{-1}$. The consequence is that at this point the heat capacity also vanishes. However, this result does not directly indicate a breakdown of thermodynamics since we only allow decay channels that change the black hole charge-to-mass ratio in a constrained manner. In the canonical ensemble (fixed charge), a vanishing heat capacity reflects that the change in mass becomes independent of the temperature. Schwinger pair production on the other hand, does not depend directly on temperature. A vanishing heat capacity at fixed $z$ therefore does not reflect a similar breakdown of thermodynamics.
This different behavior is reflected in Fig. 5 where we can tune the charge-to-mass ratio to result in a zero in the heat capacity. While the heat capacity at fixed charge is always positive along the lukewarm line (away from the Nariai limit), allowing for charge exchange we see that the heat capacity at fixed charge-to-mass ratio can have both signs.
One intriguing aspect of working at fixed $z$ is that it allows one to tune the approach to the intersection of the lukewarm and Nariai line. As already mentioned, working at fixed charge this point has a diverging heat capacity suggestive of a phase transition when crossing the lukewarm line. By tuning $z$ to obtain a finite value of the heat capacity, it is perhaps possible to better characterize the properties of this transition.
4 Near-Extremal Expansion
Before we conclude, in this section we will give a complete comparison of the different thermodynamic quantities using the $\gamma=1$ and Bousso-Hawking normalization. We will focus on the different near-extremal limits ((charged) Nariai, cold) in an ensemble of fixed charge and tabulate the behavior of various thermodynamic quantities in Tab. 3. The ultracold case will be treated separately.
The general approach we adopt is to expand the different horizon radii and $r_{\cal O}$ around their near-extremal values.
$r_{a,b,c,\mathcal{O}}=\sum_{i=0}r_{ai,bi,ci,\mathcal{O}i}\epsilon^{i}\,$~, (4.1)
where $0\leq\epsilon\ll 1$ is a dimensionless and positive expansion parameter. This approach is complementary to our previous results and will therefore also serve as a consistency check.
A subset of the coefficients in this expansion will be fixed by requiring variations of $\ell_{4}$ and the charge $Q$ to vanish, in addition to imposing $f^{\prime}(r_{\cal O})=0$ up to the relevant power in $\epsilon$. We summarize the relevant findings per case below.
Uncharged Nariai Black Hole.
Here we take $r_{\mathcal{O}0}=r_{b0}=r_{c0}=r_{0}$ and $r_{ai}=0$ for all $i\geq 0$. Explicitly this ansatz becomes
$r_{a}=0\,,\quad r_{b}=r_{0}-r_{0}\epsilon+r_{b2}\epsilon^{2}+r_{b3}\epsilon^{3}\,$, (4.2)
$r_{c}=r_{0}+r_{c1}\epsilon+r_{c2}\epsilon^{2}+r_{c3}\epsilon^{3}\,,\quad r_{\mathcal{O}}=r_{0}+r_{\mathcal{O}1}\epsilon+r_{\mathcal{O}2}\epsilon^{2}+r_{\mathcal{O}3}\epsilon^{3}\,$, (4.3)
where, without loss of generality, we used the rescaling freedom of $\epsilon$ and the other coefficients to put $r_{b1}=-r_{0}$. This ensures that $\epsilon\geq 0$ and dimensionless, since the black hole radius decreases when moving away from the Nariai limit. We neglect terms of $\mathcal{O}(\epsilon^{4})$.
We use the constraints $f^{\prime}(r_{\mathcal{O}})=0+\mathcal{O}(\epsilon^{4})$ and $\delta\ell_{4}=\mathcal{O}(\epsilon^{4})$. We choose these to fix the coefficients of $r_{c}$ and $r_{\mathcal{O}}$, which become
$r_{c1}=r_{0}\,,\quad r_{c2}=-\frac{1}{3}(r_{0}+3r_{b2})\,,\quad r_{c3}=\frac{1}{9}(r_{0}+6r_{b2}-9r_{b3})\,$, (4.4)
$r_{\mathcal{O}1}=0\,,\quad r_{\mathcal{O}2}=-\frac{r_{0}}{2}\,,\quad r_{\mathcal{O}3}=\frac{1}{6}(r_{0}+6r_{b2})\,.$ (4.5)
The thermodynamic quantities coming from the expansions above are presented in Tab. 3, in agreement with our previous results. We emphasize that $\epsilon$ can be expressed in terms of any (dimensionless) near-extremal thermodynamic quantity.
Charged Nariai Black Holes.
Akin to the uncharged Nariai case, we take $r_{\mathcal{O}0}=r_{b0}=r_{c0}=r_{0}$. The fact that we allow for charge translates to allowing for a non-vanishing value for $r_{a}$. Explicitly, this ansatz becomes
$r_{a}=r_{a0}+r_{a1}\epsilon+r_{a2}\epsilon^{2}+r_{a3}\epsilon^{3}\,,\quad r_{b}=r_{0}-r_{0}\epsilon+r_{b2}\epsilon^{2}+r_{b3}\epsilon^{3}\,$, (4.6)
$r_{c}=r_{0}+r_{c1}\epsilon+r_{c2}\epsilon^{2}+r_{c3}\epsilon^{3}\,,\quad r_{\mathcal{O}}=r_{0}+r_{\mathcal{O}1}\epsilon+r_{\mathcal{O}2}\epsilon^{2}+r_{\mathcal{O}3}\epsilon^{3}\,$, (4.7)
where again rescaled $r_{b1}$ without any loss of generality and consistent with $\epsilon\geq 0$ and dimensionless. On top of requiring $f^{\prime}(r_{\mathcal{O}})=0+\mathcal{O}(\epsilon^{4})$ and $\delta\ell_{4}=\mathcal{O}(\epsilon^{4})$ to fix the coefficients of $r_{\mathcal{O}}$ and $r_{c}$, we impose $\delta Q=\mathcal{O}(\epsilon^{4})$ in order to fix the coefficients of $r_{a}$ and enforce the canonical ensemble. The explicit expressions for these coefficients are straightforward to obtain, but large and not very insightful. We will therefore not present them here. [^9] We are happy to provide them upon request. After using the expansions above as input we obtain the thermodynamic quantities as presented in Tab. 3.
Note that $3r_{0}^{2}\leq\ell^{2}_{4}\leq 6r_{0}^{2}$, where the lower (upper) limit corresponds to the uncharged Nariai (ultracold) limit. The value $4r_{0}^{2}=\ell^{2}_{4}$ corresponds to the intersection of the lukewarm and Nariai line at which point $\tilde{C}_{b}^{Q}$ diverges, which we already discussed before. The different heat capacities agree with our previous results.
Cold Black Holes.
For cold black holes, at leading order in $\epsilon$ we find that $f(r_{\cal O})$ is constant, rather than $\epsilon^{-1}$. This means that $C^{Q}_{b}$ and $\tilde{C}^{Q}_{b}$ will be identical to leading order in $\epsilon$, both going to zero as $\epsilon\to 0$. Thus, a possible breakdown of thermodynamics can also be determined using $C^{Q}_{b}$ which has been computed previously in [10, 28].
The expansion takes the form
$r_{a}=r_{0}+r_{a1}\epsilon+r_{a2}\epsilon^{2}+r_{a3}\epsilon^{3}\,,\quad r_{b}=r_{0}+r_{0}\epsilon+r_{b2}\epsilon^{2}+r_{b3}\epsilon^{3}\,$, (4.8)
$r_{c}=r_{c0}+r_{c1}\epsilon+r_{c2}\epsilon^{2}+r_{c3}\epsilon^{3}\,,\quad r_{\mathcal{O}}=r_{\mathcal{O}0}+r_{\mathcal{O}1}\epsilon+r_{\mathcal{O}2}\epsilon^{2}+r_{\mathcal{O}3}\epsilon^{3}\,.$ (4.9)
Note that we now took $r_{b1}=r_{0}$ since the horizon radius of cold black holes increases when moving away from extremality. We note that $\ell_{4}^{2}\geq 6r_{0}^{2}$.
Again, the coefficients are fixed by imposing $f^{\prime}(r_{\mathcal{O}})=0+\mathcal{O}(\epsilon^{4})$, $\delta\ell_{4}=\mathcal{O}(\epsilon^{4})$ and $\delta Q=\mathcal{O}(\epsilon^{4})$. As for the charged Nariai case, the explicit expressions are rather long and we omit them here. The location of $r_{{\cal O}0}$ can be determined via the equation
$\ell_{4}^{2}=\frac{r_{{\cal O}0}^{3}}{r_{0}}+r_{0}r_{{\cal O}0}+3r_{0}^{2}+r_{{\cal O}0}^{2}~.$ (4.10)
The results are displayed in Tab. 3, but we denote a few of the more lengthy quantities separately here.
$\tilde{T}_{b}^{\rm cold}$ $=\frac{\epsilon(\ell_{4}^{2}-6r_{0}^{2})}{2\pi\ell_{4}^{2}r_{0}}\sqrt{\frac{r_{{\cal O}0}r_{0}\ell_{4}^{2}}{(r_{{\cal O}0}-r_{0})^{3}(r_{{\cal O}0}+r_{0})}}+{\cal O}(\epsilon^{2})~,$ (4.11)
$\tilde{T}_{c}^{\rm cold}$ $=\frac{\left(\sqrt{\ell_{4}^{2}-2r_{0}^{2}}-2r_{0}\right)\left(\ell_{4}^{2}-2r_{0}\left(\sqrt{\ell_{4}^{2}-2r_{0}^{2}}+r_{0}\right)\right)}{2\pi\ell_{4}\left(r_{0}-\sqrt{\ell_{4}^{2}-2r_{0}^{2}}\right){}^{2}}\sqrt{\frac{r_{0}r_{{\cal O}0}}{(r_{{\cal O}0}-r_{0})^{3}(r_{0}+r_{{\cal O}0})}}+{\cal O}(\epsilon)~,$
$\tilde{\Phi}_{b}^{\rm cold}$ $=\frac{\sqrt{r_{0}^{2}-\frac{3r_{0}^{4}}{\ell_{4}^{2}}}\left(r_{{\cal O}0}(1-\epsilon)-r_{0}\right)}{2\sqrt{\pi G_{4}}r_{0}r_{{\cal O}0}}\sqrt{\frac{\ell_{4}^{2}r_{0}r_{{\cal O}0}}{\left(r_{{\cal O}0}-r_{0}\right){}^{3}\left(r_{0}+r_{{\cal O}0}\right)}}+{\cal O}(\epsilon^{2})~,$
$\tilde{\Phi}_{c}^{\rm cold}$ $=-\frac{r_{0}\left(r_{{\cal O}0}\left(r_{0}+r_{{\cal O}0}\right)-\sqrt{r_{0}\left(r_{0}+r_{{\cal O}0}\right)\left(r_{0}^{2}+r_{{\cal O}0}^{2}\right)}\right)}{2\sqrt{\pi G_{4}\left(r_{{\cal O}0}-r_{0}\right){}^{3}\left(r_{0}+r_{{\cal O}0}\right)\left(r_{0}r_{{\cal O}0}+r_{0}^{2}+r_{{\cal O}0}^{2}\right)}}+{\cal O}(\epsilon)~.$
Ultracold Black Hole.
The ultracold case is more subtle because, as explained, it is neccesary to consider trajectories that obey $\delta Q/\delta M=2\sqrt{2\pi G_{4}}$. Instead of following the previous approach of fixing coefficients in a general fashion we note that an expansion of the form
$r_{b}=\frac{\ell_{4}}{\sqrt{6}}-w_{2}\epsilon~,\quad r_{c}=\frac{\ell_{4}}{\sqrt{6}}+(w_{1}+w_{2})\epsilon~,\quad r_{\cal O}=\frac{\ell_{4}}{\sqrt{6}}+\sqrt{\frac{w_{1}^{2}+w_{1}w_{1}+w_{2}^{2}}{3}}~.$ (4.12)
solves the constraints $f^{\prime}(r_{\cal O})=0+{\cal O}(\epsilon^{3})$ and $\delta Q/\delta M=2\sqrt{2\pi G_{4}}+{\cal O}(\epsilon^{3})$.
Using these expansions, all relevant thermodynamic quantities can be obtained in a straightforward manner. Instead of giving a full list, we just highlight the most important results. Due the fact that $\sqrt{f(r_{\cal O})}={\cal O}(\epsilon^{3/2})$ we find that the temperatures have a different scaling with $\epsilon$. For $\gamma=1$ normalization we have $T_{b,c}\sim\epsilon^{2}$ whereas for Bousso-Hawking normalization $\tilde{T}_{b,c}\sim\epsilon^{1/2}$. Furthermore, the chemical potentials with Bousso-Hawking normalization diverge in the ultracold limit (the (physical) electric field remains finite). Lastly, we note that the near-extremal value of the heat capacity is equal in both normalizations and given by
$C_{b}^{z}=\tilde{C}_{b}^{z}=-\sqrt{\frac{2}{3}}\frac{\pi\ell_{4}w_{2}\epsilon}{G_{4}}+{\cal O}(\epsilon^{2})~,$ (4.13)
where $z=2\sqrt{2\pi G_{4}}$.
Uncharged Nariai Charged Nariai Cold
$S_{b}$ $\frac{\pi r_{0}^{2}}{G_{4}}\left(1-2\epsilon\right)$ $\frac{\pi r_{0}^{2}}{G_{4}}\left(1-2\epsilon\right)$ $\frac{\pi r_{0}^{2}}{G_{4}}\left(1+2\epsilon\right)$
$S_{c}$ $\frac{\pi r_{0}^{2}}{G_{4}}\left(1+2\epsilon\right)$ $\frac{\pi r_{0}^{2}}{G_{4}}\left(1+2\epsilon\right)$ $\frac{\pi\left(\ell_{4}^{2}-r_{0}\left(2\sqrt{\ell_{4}^{2}-2r_{0}^{2}}+r_{0}\right)\right)}{G_{4}}$
$T_{b}$ $\frac{\epsilon}{2\pi r_{0}}$ $\frac{\epsilon}{2\pi r_{0}}\left(\frac{6r_{0}^{2}}{\ell_{4}^{2}}-1\right)$ $\frac{\epsilon}{2\pi r_{0}}\left(1-\frac{6r_{0}^{2}}{\ell_{4}^{2}}\right)$
$T_{c}$ $\frac{\epsilon}{2\pi r_{0}}$ $\frac{\epsilon}{2\pi r_{0}}\left(\frac{6r_{0}^{2}}{\ell_{4}^{2}}-1\right)$ $\frac{\sqrt{\ell_{4}^{2}-2r_{0}^{2}}\left(\sqrt{\ell_{4}^{2}-2r_{0}^{2}}-2r_{0}\right)^{2}}{2\pi\ell_{4}^{2}\left(r_{0}-\sqrt{\ell_{4}^{2}-2r_{0}^{2}}\right)^{2}}$
$\tilde{T}_{b}$ $\frac{1}{2\pi r_{0}}\left(1+\frac{2}{3}\epsilon\right)$ $\frac{\sqrt{6r_{0}^{2}-\ell_{4}^{2}}}{2\pi\ell_{4}r_{0}}\left(1+\frac{2\epsilon\left(\ell_{4}^{2}-4r_{0}^{2}\right)}{\ell_{4}^{2}-6r_{0}^{2}}\right)$ $\tilde{T}_{b}^{\rm cold}$
$\tilde{T}_{c}$ $\frac{1}{2\pi r_{0}}\left(1-\frac{2}{3}\epsilon\right)$ $\frac{\sqrt{6r_{0}^{2}-\ell_{4}^{2}}}{2\pi\ell_{4}r_{0}}\left(1-\frac{2\epsilon\left(\ell_{4}^{2}-4r_{0}^{2}\right)}{\ell_{4}^{2}-6r_{0}^{2}}\right)$ $\tilde{T}_{c}^{\rm cold}$
$\Phi_{b}$ 0 $\sqrt{\frac{\ell_{4}^{2}-3r_{0}^{2}}{4\pi G_{4}\ell_{4}^{2}}}(1+\epsilon)$ $\sqrt{\frac{r_{0}\ell_{4}^{2}-3r_{0}^{3}}{4\pi G_{4}r_{0}\ell_{4}^{2}}}(1-\epsilon)$
$\Phi_{c}$ 0 $\sqrt{\frac{\ell_{4}^{2}-3r_{0}^{2}}{4\pi G_{4}\ell_{4}^{2}}}(1-\epsilon)$ $\frac{\sqrt{r_{0}^{2}\ell_{4}^{2}-3r_{0}^{4}}}{\sqrt{4\pi G_{4}\ell_{4}^{2}}\left(\sqrt{\ell_{4}^{2}-2r_{0}^{2}}-r_{0}\right)}$
$\tilde{\Phi}_{b}$ 0 $\sqrt{\frac{\ell_{4}^{2}-3r_{0}^{2}}{4\pi G_{4}(6r_{0}^{2}-\ell_{4}^{2})}}\left(1-\frac{2r_{0}^{2}\epsilon}{\ell_{4}^{2}-6r_{0}^{2}}\right)$ $\tilde{\Phi}_{b}^{\rm cold}$
$\tilde{\Phi}_{c}$ 0 $\sqrt{\frac{\ell_{4}^{2}-3r_{0}^{2}}{4\pi G_{4}(6r_{0}^{2}-\ell_{4}^{2})}}\left(-1-\frac{2r_{0}^{2}\epsilon}{\ell_{4}^{2}-6r_{0}^{2}}\right)$ $\tilde{\Phi}_{c}^{\rm cold}$
$Q^{2}$ 0 $\frac{4\pi r_{0}^{2}\left(\ell_{4}^{2}-3r_{0}^{2}\right)}{G_{4}\ell_{4}^{2}}$ $\frac{4\pi r_{0}^{2}\left(\ell_{4}^{2}-3r_{0}^{2}\right)}{G_{4}\ell_{4}^{2}}$
$\ell_{4}^{2}$ $3r_{0}^{2}$ $3r_{0}^{2}+2r_{0}r_{a0}+r_{a0}^{2}$ $3r_{0}^{2}+2r_{0}r_{c0}+r_{c0}^{2}$
$C_{b}^{Q}$ $\frac{2\pi r_{0}^{2}\epsilon}{G_{4}}$ $\frac{2\pi r_{0}^{2}\epsilon}{G_{4}}$ $\frac{2\pi r_{0}^{2}\epsilon}{G_{4}}$
$\tilde{C}_{b}^{Q}$ $\frac{3\pi r_{0}^{2}}{G_{4}}\left(1-\frac{11}{6}\epsilon\right)$ $\frac{\pi r_{0}^{2}\left(6r_{0}^{2}-\ell_{4}^{2}\right)}{G_{4}\left(4r_{0}^{2}-\ell_{4}^{2}\right)}-\frac{\pi r_{0}\epsilon\left(49\ell_{4}^{2}r_{0}^{3}-4\ell_{4}^{4}r_{0}-122r_{0}^{5}\right)}{2G_{4}\left(\ell_{4}^{2}-4r_{0}^{2}\right)^{2}}$ $\frac{2\pi r_{0}^{2}\epsilon}{G_{4}}$
Table 3: Near-extremal expansion of the different thermodynamic quantities for (un)charged Nariai and cold black holes where we ignored terms of ${\cal O}(\epsilon^{2})$ and smaller. Some of the more lengthy expressions are presented in (4.11). The results for the heat capacity agree with our previous results. The expansion parameter $\epsilon$ is dimensionless and positive. For the electric potential we took the gauge $\lambda=0$ for $\gamma=1$ and $\tilde{\lambda}=r_{\cal O}^{-1}$ for Bousso-Hawking normalization.
Breakdown of the Semi-Classical Approximation.
From the above results, we can deduce where the thermodynamic description breaks down due to the (absolute value of the) heat capacity becoming smaller than one. Whenever this happens, one might expect large (quantum) corrections to modify the behavior of the partition function, as is the case for log-$T$ corrections.
For cold black holes, the normalization of the Killing vector does not influence the value of the heat capacity to leading order in $\epsilon$. The heat capacity vanishes, the thermodynamic description breaks down and we expect log-$T$ corrections to the partition function.
For Nariai black holes, we found that sufficiently far away from the ultracold point the (absolute value of the) heat capacity remains large when using Bousso-Hawking normalization. This signals no breakdown of the thermodynamic description, therefore questioning the presence of large (quantum) corrections. This should be contrasted with $\gamma=1$ normalization where the heat capacity is zero along the entire Nariai branch. It would therefore be of great interest to compute the one-loop determinant of near-Nariai black holes using Bousso-Hawking normalization to see if this intuition is correct and whether the normalization of the Killing vector plays an important role. Existing results on one-loop determinants for de Sitter black holes include [34, 35, 36, 37, 28, 38, 39] and some of these references did report large quantum corrections in the Nariai limit (see e.g. [36, 37, 28] for recent results). These results, however, seem to depend on a complexification of the geometry describing the Milne patch of the near-horizon de Sitter region, instead of the static patch that we focus on. Therefore, these one-loop results apply in a different regime. Nonetheless, it would be interesting to better understand the relation between these computations and our thermodynamic approach.
Lastly, for ultracold black holes the value of the heat capacity is again equal for both choices of normalization so we do expect a breakdown of the thermodynamic decscription and, most likely, large corrections. Explicitly computing these corrections in the ultracold limit, however, has remained a challenge.
5 Conclusions
In this work, we investigated the thermodynamics of four-dimensional electrically charged de Sitter black holes, focusing on the ambiguities that arise from the choice of normalization of the timelike Killing vector. We demonstrated that the conventional Gibbons-Hawking normalization leads to an uncommon interpretation of the black hole temperature; it expresses the surface gravity at the horizon as an acceleration measured in the frame of an accelerating observer that is located behind a horizon.
To remedy this, we adopted the Bousso-Hawking normalization of the Killing vector. This choice expresses the different thermodynamic quantities with respect to the frame of the unique freely falling observer at a fixed radial location in between the (outer) black hole horizon and cosmological horizon. This provides a more physically coherent picture, and we derived the corresponding first laws of thermodynamics for both the black hole and cosmological horizon. This led to the introduction of a new mass parameter $\tilde{M}$ whose variation equals the variation of the mass parameter $M$ (defined in the normalization used by Gibbons and Hawking) multiplied by a redshift factor that adapts it to the frame of the freely falling observer.
Our primary result is the calculation of the black hole heat capacity using the Bousso-Hawking normalization, with a focus on its behavior near the three extremal limits: cold, Nariai, and ultracold. As established in the literature and discussed in the main text, a heat capacity that approaches order unity at low temperatures typically signals the breakdown of the semi-classical thermodynamic description. For asymptotically flat black holes, this phenomenon has been understood to stem from quantum corrections to the Euclidean path integral’s saddle-point approximation. While these corrections are negligible at high temperatures, they dominate at low temperatures, yielding characteristic logarithmic (log-$T$) corrections to the partition function.
In contrast to results for the heat capacity using the normalization by Gibbons and Hawking, we find that the heat capacity remains finite and large for Nariai black holes, sufficiently far away from the ultracold point. For cold and ultracold black holes, however, we find the same result for the heat capacity, which vanishes in these extremal limits. Our results therefore suggest that there is no breakdown of the thermodynamic description along the Nariai branch (away from the ultracold point).
For this reason, we also do not expect to see large (quantum) corrections in this limit, and it would be very interesting to confirm this prediction by an explicit one-loop computation that takes into account the physical normalization of the Killing vector. Intriguingly, the results of [38] on Kerr-de Sitter black holes seem to point into this direction. Using the DHS formula [40], the authors of [38] identified whether log-$T$ corrections are present in the different extremal limits from the quasi-normal mode spectrum. Similar to our results, they find log-$T$ corrections for the cold and ultracold branch, but an absence of log-$T$ corrections for Nariai black holes. Although their approach focuses on rotating black holes and does not seem to rely on Bousso-Hawking normalization, we find this connection intriguing and we believe it warrants further investigation. [^10] We believe that their approach effectively implements our choice of normalization through the boundary conditions of the quasinormal modes.
It is also instructive to compare our findings with those of [36, 28], which reported log-$T$ corrections for both cold and charged Nariai black holes. These studies derived corrections to the partition function by analyzing the (near-) zero modes of the Euclidean gravitational path integral. Specifically, in the Nariai limit, these zero modes are related to what is sometimes referred to as the large diffeomorphisms of the near-horizon de Sitter region. Since the appropriate evaluation of these modes involves a complexification of the geometry that takes one outside of the static patch, this seems to implicitly rely on a choice of reference frame different from ours. We therefore expect that their approach, phrased in terms of the heat capacity, corresponds to a normalization of the Killing vector different from the Bousso-Hawking one.
In summary, to fully understand the relation between the different approaches that study corrections to near-extremal black holes in de Sitter space, it is necessary to explicitly compute the one-loop quantum corrections and understand the importance of a physical observer. This perspective seems to be in line with other recent developments that stress the role of the observer in de Sitter space [41, 42, 43, 44, 45, 21].
Acknowledgements
We would like to thank Manus Visser for fruitful discussions at an early stage of this project. LA is supported by the President’s Postdoctoral Fellowship Program at the University of Minnesota and the Heising-Simons Foundation under the “Observational Signatures of Quantum Gravity” collaboration grant 2021-2818. LA thanks the organizers and participants of the workshop “Quantum Gravity in Duluth” where part of this work was presented and acknowledges the University of Wisconsin-Madison for hospitality. PL and GS are supported in part by the DOE grant DE-SC0017647. The work of WS is supported by Starting Grant 2023-03373 from the Swedish Research Council.
Appendix A Quartic Polynomials
Here we give the general solution for the roots of a quartic polynomial. Consider a quartic polynomial given by
$P_{4}=ar^{4}+br^{3}+cr^{2}+dr+e~.$ (A.1)
The four roots $P_{4}=0$ are given by [46]
$r_{1,2}$ $=-\frac{b}{4a}-S\pm\frac{1}{2}\sqrt{-4S^{2}-2p+\frac{q}{S}}~,$ (A.2)
$r_{3,4}$ $=-\frac{b}{4a}+S\pm\frac{1}{2}\sqrt{-4S^{2}-2p-\frac{q}{S}}~,$
with
$p$ $=\frac{8ac-3b^{2}}{8a^{2}}~,$ (A.3)
$q$ $=\frac{b^{3}-4abc+8a^{2}d}{8a^{3}}~,$
and
$S$ $=\frac{1}{2}\sqrt{-\frac{2}{3}p+\frac{1}{3a}\left(R+\frac{\Delta_{0}}{R}\right)}~,$ (A.4)
$R$ $=\left(\frac{\Delta_{1}+\sqrt{\Delta_{1}^{2}-4\Delta_{0}^{3}}}{2}\right)^{1/3}~,$
where
[…]
Appendix B Derivation of Thermodynamic Quantities
In this appendix, we specify the geometric setup for computing various thermodynamic quantities.
B.1 General Expressions
We consider a $d$-dimensional spacetime and define a spacelike co-dimension 1 hypersurface $\Sigma$ with timelike unit normal $n_{\mu}$. We also consider a co-dimension 1 timelike surface $B$ with spacelike unit normal $r_{\mu}$. From the induced metric on $B$, we can construct the extrinsic curvature $k^{\mu\nu}$ with trace $k$. We use $\sigma_{\mu\nu}$ to denote the induced metric on the intersection $\Sigma\cap B$.
Following [12], for a timelike Killing vector $\xi^{\mu}$, we define the local acceleration associated with this trajectory by
$a_{\mu}=\frac{\nabla_{\mu}(\sqrt{-\xi^{\nu}\xi_{\nu}})}{\sqrt{-\xi^{\rho}\xi_{\rho}}}~.$ (B.1)
Interpreting $B$ as a York boundary, we define $s^{\mu\nu}$ to be the spatial stress contribution to the Brown-York stress tensor [47]
$s^{\mu\nu}=\frac{1}{8\pi G_{d}}\left(-k^{\mu\nu}+(r_{\rho}a^{\rho}+k)\sigma^{\mu\nu}\right)~.$ (B.2)
Using this definition, we can define the pressure associated to the York boundary as
$P_{B}=\frac{1}{d-2}s^{\mu\nu}\sigma_{\mu\nu}~.$ (B.3)
Another quantity that we need to introduce is the anti-symmetric Killing potential $\omega^{\mu\nu}$ that satisfies
$\nabla_{\nu}\omega^{\nu\mu}=\xi^{\mu}~.$ (B.4)
Given a Killing vector, the Killing potential is non unique and can be modified by a term $\omega^{\mu\nu}\to\omega^{\nu}+\lambda^{\mu\nu}$, where $\lambda^{\mu\nu}$ is an anti-symmetric tensor that satisfies $\nabla_{\nu}\lambda^{\nu\mu}=0$. From the Killing potential, one can define the thermodynamic volume between two different timelike surfaces $B_{1}$ and $B_{2}$. For example, for $B_{b}$ the black hole horizon and $B_{\infty}$ infinity the thermodynamic volume is given by
$\int_{\Sigma\cap B_{\infty}}\mathrm{d}A\,r_{\mu}n_{\nu}\left(\omega^{\mu\nu}-\omega^{\mu\nu}_{0}\right)-\int_{\Sigma\cap B_{b}}\mathrm{d}A\,r_{\mu}n_{\nu}\omega^{\mu\nu}~,$ (B.5)
where we included the possibility of adding a counter term $\omega^{\mu\nu}$ to subtract of a divergence. $\mathrm{d}A$ is the area element on the intersection $\Sigma\cap B$.
B.2 Charged de Sitter Black Holes
Now let us evaluate these quantities explicitly for four-dimensional charged de Sitter black holes with Killing vector $\xi=\gamma\partial_{t}$. We find that the pressure at a location of constant $r$ is given by
$P_{r}^{-1}=8\pi G_{4}\ell_{4}^{2}r^{2}\sqrt{-\frac{\left(r-r_{b}\right)\left(r-r_{c}\right)\left(r_{b}\left(r_{c}+r\right)+r_{b}^{2}+r_{c}\left(r_{c}+r\right)-\ell_{4}^{2}+r^{2}\right)}{\ell_{4}^{2}\left(r_{b}r_{c}\left(r_{b}r_{c}+r_{b}^{2}+r_{c}^{2}-\ell_{4}^{2}\right)+r^{2}\left(\ell_{4}^{2}-3r^{2}\right)\right){}^{2}}}$ (B.6)
To derive the Killing potential, we take the ansatz that the only non-zero components are $\omega^{tr}=-\omega^{rt}$ and we assume the components only depend on the radial coordinate. We then find
$\omega^{tr}=\gamma\left(-\frac{r}{3}+\alpha\frac{r_{b}^{3}}{r^{2}}\right)~,$ (B.7)
with $\alpha$ a constant that captures the ambiguity in the Killing potential. To compute the thermodynamic volume between the black hole horizon and infinity, we need to specify a counter term due to a divergence. We specify $\omega^{\mu\nu}_{0}$ to be the Killing potential for empty de Sitter space, which suggests us to set $\alpha=0$ to remove the divergence at $r=0$. Doing so, we find that (B.5) evaluates to
$V_{b}=\gamma\frac{4}{3}\pi r_{b}^{3}~.$ (B.8)
If we compute the thermodynamic volume between two surfaces $B_{1}$ and $B_{2}$ at (finite) radius $r=r_{1}$ and $r=r_{2}$ we find that there is no need to regularize, $\alpha$ drops out and we are left with
$V_{r_{2}}-V_{r_{1}}=\gamma\frac{4}{3}\pi(r_{2}^{3}-r_{1}^{3})~.$ (B.9)